### v1

# Path Eccentricity and Forbidden Induced Subgraphs

Abstract
[…]
1 Introduction
A Hamiltonian path in a graph $G$ is a path in $G$ whose vertex set is $V(G)$; a Hamiltonian cycle is defined similarly. The problems of determining whether a given graph has a Hamiltonian path or a Hamiltonian cycle are both NP-complete [GJ79, Kar72], which motivates the search for sufficient conditions for their existence (see, e.g., the survey [Gou14]). One of the early results in the area is the following result of Duffus, Jacobson, and Gould [DJG81] (see also [She91, BD03, Kel06]).
Theorem 1.1.
[…]
Here, a claw is the complete bipartite graph $K_{1,3}$, the net is the graph obtained by appending a pendant edge to each vertex of the complete graph $K_{3}$, and, given a set $\mathcal{F}$ of graphs, a graph $G$ is said to be $\mathcal{F}$-free if $G$ does not have an induced subgraph isomorphic to any member of $\mathcal{F}$. If =$\mathcal{F}=\{F\}$ for some graph $F$, we also say that $G$ is $F$-free.
The result of Theorem 1.1 is tight, since the claw and the net do not have a Hamiltonian path. Let us remark that, in general, if $G$ is a graph that has a Hamiltonian path and $H$ is a connected induced subgraph of $G$, then $H$ may fail to have a Hamiltonian path. For example, if $H$ is any $n$-vertex graph and $G$ is a graph obtained from $H$ by adding to it $n$ new vertices fully adjacent to the vertices of $H$, then $G$ has a Hamiltonian path, even though $H$ may fail to have one. In particular, this construction could be applied to the the claw and the net.
If we instead require that all connected induced subgraph have a Hamiltonian path, Theorem 1.1 implies the following characterization.
Corollary 1.2.
For every graph $G$, the following statements are equivalent:
  • Every connected induced subgraph of $G$ has a Hamiltonian path.
[…]
In 2023, Gómez and Gutiérrez [GG23] generalized the concept of Hamiltonian paths by defining the notion of path eccentricity of graphs. Given a connected graph $G$ and a path $P$ in $G$, the eccentricity of $P$ is the maximum distance from a vertex in $G$ to a vertex in $P$. Hence, in particular, a path $P$ has eccentricity 00 if and only if $P$ is a Hamiltonian path, and $P$ has eccentricity at most $1$ if and only if $P$ is dominating, that is, every vertex not in the path has a neighbor on the path. More generally, given an integer $k\geqslant 0$, a path $P$ is said to be $k$-dominating if it has eccentricity at most $k$. The path eccentricity of a connected graph $G$ is denoted by $\operatorname{\mathbf{pe}}(G)$ and defined as the minimum eccentricity of a path in $G$. Hence, $\operatorname{\mathbf{pe}}(G)=0$ if and only if $G$ has a Hamiltonian path, and $\operatorname{\mathbf{pe}}(G)\leqslant 1$ if and only if $G$ has a dominating path. It is not difficult to see that a graph $G$ has a dominating path if and only if $G$ has a spanning caterpillar, that is, a spanning subgraph that is a caterpillar (a tree possessing a dominating path). Hence, understanding graphs that admit dominating paths is motivated by their applications to the graph burning and a variant of the cops and robber game. More precisely:
  • The burning graph conjecture (see Bonato, Janssen, and Roshanbin [BJR16]) is known to be true for graphs admitting a spanning caterpillar.
  • If a graph $G$ admits a spanning caterpillar $T$, then the minimum number of cops sufficient for the cop player to win on the graph $G$ against the robber moving at the speed of $s$ is at most $ps$ if $G$ is a subgraph of the graph obtained from $T$ by adding an edge between each pair of vertices that are at distance at most $p$ in $T$ (see Fomin et al. [FGK+10, Lemma 4]).
Since adding a universal vertex to any connected graph $H$ results in a graph $G$ with path eccentricity at most $1$, path eccentricity is not monotone under vertex deletion. Nevertheless, and in line with the above discussion preceding Corollary 1.2, which characterizes graphs each connected induced subgraph of which has path eccentricity 00, analogous questions can be addressed for higher values of path eccentricity.
Question 1.3.
For a positive integer $k$, what are the graphs $G$ such that each connected induced subgraph of $G$ has path eccentricity at most $k$?
For each fixed $k\geqslant 1$, there are two obstructions to path eccentricity less than $k$ that can be obtained from the claw and the net, as follows. Given a positive integer $k$, let us denote by $S_{k}$ the $k$-subdivided claw, that is, the graph obtained from the claw by replacing each of its edges with a path of length $k$ (in particular, $S_{1}$ is the claw). Furthermore, we denote by $T_{k}$ the $k$-subdivided net, that is, the graph obtained by appending a path of length $k$ to each vertex of the complete graph $K_{3}$ (in particular, $T_{1}$ is the net). See Figure 1 for an example.
Figure 1: Visualization of graphs $S_{2}$ and $T_{2}$.
It is easy to observe that $\operatorname{\mathbf{pe}}(S_{k})=\operatorname{\mathbf{pe}}(T_{k})=k$ for each positive integer $k$. Consequently, if $G$ is a graph such that each connected induced subgraph of $G$ has path eccentricity less than $k$, then $G$ is $\{S_{k},T_{k}\}$-free. As the main result of this paper, we show that this necessary condition is in fact also sufficient. We thus obtain the following unifying theorem, valid for all positive integers $k$ (and thus capturing also Corollary 1.2, which corresponds to the case $k=1$).
Theorem 1.4.
For every integer $k\geqslant 1$ and every graph $G$, the following statements are equivalent:
  • Every connected induced subgraph $H$ of $G$ satisfies $\operatorname{\mathbf{pe}}(H)<k$.
  • $G$ is $\{S_{k},T_{k}\}$-free.
The case $k=2$ of Theorem 1.4 characterizes graphs every connected induced subgraph of which has a dominating path, and implies the following.
Corollary 1.5.
Every connected $\{S_{2},T_{2}\}$-free graph has a dominating path.
Corollary 1.5 unifies and generalizes several results from the literature regarding sufficient conditions for the existence of a dominating path:
  1. 1.
    One of the oldest results along these lines is the result of Bacsó and Tuza in 1990 (see [BT90]) stating that every connected $P_{5}$-free graph has a dominating clique or a dominating $P_{3}$ (we denote by $P_{k}$ the $k$-vertex path). It follows that every connected $P_{5}$-free graph has a dominating path.
  2. 2.
    A graph $G$ is said to be AT-free if it does not have an independent set $I$ such that $|I|=3$ and for every vertex $v\in I$, the remaining two vertices in $I$ are in the same component of $G-N[v]$. Every connected AT-free graph has a dominating path (see Corneil, Olariu, and Stewart [COS97], as well as [COS95]).
  3. 3.
    A graph $G$ is biconvex if it is bipartite, with parts $A$ and $B$ that can each be linearly ordered so that for each vertex $v$ of $G$, the neighborhood of $v$ in the part not containing $v$ forms a consecutive segment of vertices with respect to the linear ordering. Gómez and Gutiérrez [GG23] and independently Antony et al. [ADG+24] proved that every connected biconvex graph has a dominating path (or, equivalently, a spanning caterpillar).
Indeed, each of the previously stated cases corresponds to a subclass of the class of $\{S_{2},T_{2}\}$-free graphs, hence, Corollary 1.5 implies all these results.
Further consequences of Theorem 1.4 are related to a recent work of Bastide, Hilaire, and Robinson [BHR25]. The authors explored a relation between the path eccentricity of a graph and a certain property of its adjacency matrix (or equivalently on the neighborhoods of its vertices). They introduced the partially augmented consecutive ones property (denoted *-C1P for short): a graph has the *-C1P if there exists an ordering of its vertices sending the open or closed neighborhood of every vertex to a consecutive set. The authors showed that graphs with the *-C1P have path eccentricity at most 2, and conjectured that they actually have path eccentricity at most 1. We show that Theorem 1.4 resolves the conjecture.
Theorem 1.6.
If a graph $G$ has the *-C1P, then $\operatorname{\mathbf{pe}}(G)\leqslant 1$.
Another result from [BHR25] is related to the following generalization of AT-freeness. A graph $G$ is said to be $k$-AT-free if $G$ does not have an independent set $I$ such that $|I|=3$ and for every vertex $v\in I$, the remaining two vertices in $I$ are in the same component of the graph $G-\{u\colon u$ is at distance at most $k$ of $v\}$. In particular, a graph is AT-free if and only if it is $1$-AT-free. In [BHR25], the authors proved that for every $k\geqslant 1$, if a graph $G$ is $k$-AT-free, then $\operatorname{\mathbf{pe}}(G)\leqslant k$. This result follows from Theorem 1.4. Indeed, it is easy to check that for every $k\geqslant 2$, the graphs $S_{k}$ and $T_{k-1}$ are not $(k-1)$-AT-free. Thus, if a graph $G$ is $(k-1)$-AT-free, then $G$ is $\{S_{k},T_{k}\}$-free, which implies by Theorem 1.4, that $\operatorname{\mathbf{pe}}(G)<k$.
For the case where a single induced subgraph is excluded, Theorem 1.4 leads to the following characterization. The symbol $+$ denotes the disjoint union of graphs, and for an integer $k\geqslant 0$ and a graph $H$, we denote by $kH$ the disjoint union of $k$ copies of $H$.
Corollary 1.7.
Let $H$ be a graph and $k\geqslant 1$. Then, the following statements are equivalent:
  • Every connected $H$-free graph $G$ has $\operatorname{\mathbf{pe}}(G)<k$.
  • $H$ is an induced subgraph of $3P_{k}$ or $P_{2k+1}+P_{k-1}$.
In the case of $3P_{2}$-free graphs, we obtain a much stronger property. We can find a path with eccentricity at most $1$ that is a longest path in the graph. While the proof of this result implies that we can transform a longest path into a longest path that is dominating in polynomial time, it is NP-hard to find a longest path in $3P_{2}$-free graphs. In fact, determining whether a given graph has a Hamiltonian path is NP-complete in the class of split graphs [Mül96] that are even $2P_{2}$-free.
Theorem 1.8.
Every connected $3P_{2}$-free graph $G$ has a longest path that is dominating. Moreover, we can find a (not necessarily longest) dominating path in $G$ in time $\mathcal{O}(n^{2}(n+m))$, where $n$ is the number of vertices and $m$ is the number of edges in $G$.
Figure 2: An example of a $\{P_{5},4P_{2}\}$-free graph no longest path of which is dominating. The vertices depicted in orange need to belong to every dominating path; however, no longest path contains the top orange vertex.
This is a very fine line, as this property is completely violated in the case of $\{S_{2},T_{2}\}$-free graphs or even $\{P_{5},4P_{2}\}$-free graphs; see Figure 2. It also cannot be generalized to higher eccentricity; see for example a $\{3P_{3},P_{7}+P_{2}\}$-free graph on Figure 3. Moreover, in $3P_{2}$-free graph there might exist a longest path that is not dominating, as depicted in Figure 4.
Figure 3: An example of a $\{3P_{3},P_{7}+P_{2}\}$-free graph no longest path of which is $2$-dominating. The vertices depicted in orange need to belong to every $2$-dominating path; however, no longest path contains the top orange vertex.
Figure 4: An example of $3P_{2}$-free graph with one of the longest paths (depicted in orange) that is not dominating.
While Theorem 1.8 actually provides us with a polynomial-time algorithm that finds a dominating path, the algorithm does not necessarily find a longest path that is dominating, unless we are given another longest path on the input.
Organization of the paper.
In Section 2, we give an overview of known results on structural domination, including a key result of Bacsó and Tuza [BT12] that we use in the proof of our main theorem (Theorem 1.4). We prove Theorems 1.4 and 1.7 in Section 3. We prove Theorem 1.6 in Section 4. We prove Theorem 1.8 in Section 5. We conclude the paper with some final remarks and open questions in Section 6.
2 Background on Structural Domination
Theorem 1.4 is closely related to the framework of structural domination. Given a class $\mathcal{D}$ of connected graphs, let us say that a graph $G$ is hereditarily dominated by $\mathcal{D}$ if each connected induced subgraph of $G$ contains a dominating set that induces a graph from $\mathcal{D}$. Bacsó, Michalak, and Tuza (see [BMT05]) gave forbidden induced subgraph characterizations of graphs hereditarily dominated by classes of complete bipartite graphs, stars, connected bipartite graphs, and complete $k$-partite graphs. Similar results were obtained by Michalak (see [Mic07]), for graphs hereditarily dominated by classes of cycles and paths, paths, trees, and trees with bounded diameter. Camby and Schaudt (see [CS16]) proved that for every $k\geqslant 4$, a graph $G$ is $P_{k}$-free if and only if it is hereditarily dominated by the class consisting of all connected $P_{k-2}$-free graphs along with the $k$-vertex cycle.
The most general result along these lines was obtained independently by Bacsó and Tuza (see [Tuz08, Bac09]). To state the result, we need one more definition. Given a connected graph $H$, the leaf graph of $H$ is the graph $H^{\text{\textleaf}}$ obtained from $H$ by attaching a leaf to each vertex $v\in V(H)$ that is not a cut-vertex (that is, the graph $H-v$ is connected). By definition, these new leaves are pairwise distinct and non-adjacent.
Theorem 2.1 (Bacsó [Bac09] and Tuza [Tuz08]).
Let $\mathcal{D}$ be a class of connected graphs that contains the class of all paths, excludes at least one connected graph, and is closed under taking connected induced subgraphs. Let $\mathcal{F}$ be the family of all connected graphs $F$ that do not belong to $\mathcal{D}$, but all proper connected induced subgraphs of $F$ belong to $\mathcal{D}$. Let $\mathcal{F}^{\text{\textleaf}}$ be the family of all leaf graphs of graphs in $\mathcal{F}$. Then, a graph $G$ is hereditarily dominated by $\mathcal{D}$ if and only if $G$ is $\mathcal{F}^{\text{\textleaf}}$-free.
The main result of [Tuz08, Bac09] is even more general, without the restriction that $\mathcal{D}$ contains the class of all paths; but since we do not need this result, we omit the precise statement.
A generalization of Theorem 2.1 to domination at distance was proved by Bacsó and Tuza (see [BT12]). A distance-$k$ dominating set in a graph $G$ is a set $D\subseteq V(G)$ such that every vertex in $G$ is at distance at most $k$ from some vertex in $D$. If, in addition, the graph $G[D]$ is connected, then $D$ is said to be a connected distance-$k$ dominating set. A graph $G$ is said to be hereditarily $k$-dominated by a class $\mathcal{D}$ of connected graphs if each connected induced subgraph of $G$ admits a distance-$k$ dominating set that induces a graph from $\mathcal{D}$. Given a connected graph $H$, the $k$-leaf graph of $H$ is the graph $H_{k}^{\text{\textleaf}}$ obtained from $H$ by attaching a pendant path of length $k$ to each vertex $v\in V(H)$ that is not a cut-vertex. Bacsó and Tuza (see [BT12]) generalized Theorem 2.1 as follows.
Theorem 2.2 (Bacsó and Tuza [BT12]).
Let $\mathcal{D}$ be a class of connected graphs that contains the class of all paths, excludes at least one connected graph, and is closed under taking connected induced subgraphs. Let $\mathcal{F}$ be the family of all connected graphs $F$ that do not belong to $\mathcal{D}$, but all proper connected induced subgraphs of $F$ belong to $\mathcal{D}$. Let $k$ be a positive integer and let $\mathcal{F}_{k}^{\text{\textleaf}}$ be the family of all $k$-leaf graphs of graphs in $\mathcal{F}$. Then, a graph $G$ is hereditarily $k$-dominated by $\mathcal{D}$ if and only if $G$ is $\mathcal{F}_{k}^{\text{\textleaf}}$-free.
Note that Theorem 2.1 is a special case of Theorem 2.2 for the case $k=1$.
3 Proof of General Theorem (Theorem 1.4)
For convenience, we restate our main theorem.
See 1.4
Proof.
Fix an integer $k\geqslant 1$. Let $G$ be a graph such that every connected induced subgraph $H$ of $G$ satisfies $\operatorname{\mathbf{pe}}(H)<k$. To show that $G$ is $\{S_{k},T_{k}\}$-free, it suffices to show that $\operatorname{\mathbf{pe}}(S_{k})\geqslant k$ and $\operatorname{\mathbf{pe}}(T_{k})\geqslant k$. Let $H$ be a graph isomorphic to either $S_{k}$ or $T_{k}$ and let $S$ be any path in $H$. We claim that $S$ has eccentricity at least $k$. Suppose not. Let $C$ be the set of vertices of degree $3$ in $H$. Then, the graph $H-V(C)$ consists of three disjoint paths $P$, $Q$, and $R$, each with $k$ vertices. Moreover, each of the paths $P$, $Q$, and $R$ contains a vertex at distance $k$ from $V(C)$. Since the eccentricity of $S$ is less than $k$, each of the paths $P$, $Q$, and $R$ contains a vertex of $S$. Since $S$ has at most two endpoints, at most two of the paths $P$, $Q$, and $R$ contain an endpoint of $P$; hence, we may assume without loss of generality that $P$ contains an internal vertex of $S$ but no endpoint of $S$. However, this implies that two distinct edges of $S$ connect $P$ with the rest of $H$, a contradiction. This shows that $\operatorname{\mathbf{pe}}(H)\geqslant k$, as claimed.
It remains to prove that if $G$ is $\{S_{k},T_{k}\}$-free, then every connected induced subgraph $H$ of $G$ has path eccentricity less than $k$, or, equivalently, that every connected $\{S_{k},T_{k}\}$-free graph has path eccentricity less than $k$. For $k=1$, the statement coincides with that of Theorem 1.1. Suppose now that $k\geqslant 2$. Let =$\mathcal{F}=\{S_{1},T_{1}\}$ and let $\mathcal{D}$ be the class of all connected $\mathcal{F}$-free graphs. Note that $\mathcal{F}_{k-1}^{\text{\textleaf}}=\{S_{k},T_{k}\}$. Consequently, since $\mathcal{D}$ satisfies the hypothesis of Theorem 2.2, it follows that every $\{S_{k},T_{k}\}$-free graph $G$ is hereditarily $(k-1)$-dominated by $\mathcal{D}$. Let $H$ be a connected induced subgraph of $G$. Then, since $G$ is hereditarily $(k-1)$-dominated by $\mathcal{D}$, there exists a distance-$(k-1)$ dominating set $D$ in $H$ such that the induced subgraph $H[D]$ belongs to $\mathcal{D}$. By Theorem 1.1, the graph $H[D]$ admits a Hamiltonian path $P$. It follows that $P$ is a path in $H$ with eccentricity less than $k$, implying that $\operatorname{\mathbf{pe}}(H)<k$. ∎
Our proof of Corollary 1.7 will be based on the following.
Lemma 3.1.
Let $k\geqslant 1$ be an integer and $H$ be a graph. Then, $H$ is an induced subgraph of both $S_{k}$ and $T_{k}$ if and only if $H$ is an induced subgraph of $3P_{k}$ or $P_{2k+1}+P_{k-1}$.
Proof.
Suppose first that $H$ is an induced subgraph of $3P_{k}$ or $P_{2k+1}+P_{k-1}$. Since $3P_{k}$ and $P_{2k+1}+P_{k-1}$ are induced subgraphs of both $S_{k}$ and $T_{k}$, it immediately follows that $H$ is an induced subgraph of both $S_{k}$ and $T_{k}$.
Suppose now that $H$ is an induced subgraph of both $S_{k}$ and $T_{k}$. If $H$ admits a vertex $u$ of degree $3$, then, since $H$ is an induced subgraph of $T_{k}$, two neighbors of $u$ are adjacent, forming a cycle. However, this is impossible, since $H$ is also a subgraph of the acyclic graph $S_{k}$. Thus $H$ has maximum degree at most $2$. This implies that $H$ is an induced subgraph of either the graph obtained by removing the vertex of degree $3$ from $S_{k}$, which is isomorphic to $3P_{k}$; or the graph obtained by removing a neighbor of the vertex of degree $3$ from $S_{k}$, which is isomorphic to $P_{2k+1}+P_{k-1}$. ∎
See 1.7
Proof.
Suppose first that $H$ is an induced subgraph of $3P_{k}$ or $P_{2k+1}+P_{k-1}$. By Lemma 3.1, $H$ is an induced subgraph of both $S_{k}$ and $T_{k}$. This implies that every connected $H$-free graph $G$ is $\{S_{k},T_{k}\}$-free and, hence, satisfies $\operatorname{\mathbf{pe}}(G)<k$ thanks to Theorem 1.4.
Suppose now that $H$ is not an induced subgraph of $3P_{k}$ or $P_{2k+1}+P_{k-1}$. Then, again by Lemma 3.1, $H$ is not an induced subgraph of either $S_{k}$ or $T_{k}$. Let $R_{k}$ be a graph in $\{S_{k},T_{k}\}$ that does not admit $H$ as induced subgraph. Then $R_{k}$ is a connected $H$-free graph and ${\operatorname{\mathbf{pe}}(R_{k})\geqslant k}$ (by Theorem 1.4). ∎
4 Implications for the *-C1P Property (Theorem 1.6)
We start with the following lemma, which is an unpublished result of Bastide, Robinson, and the second author, obtained during the writing of [BHR25]. We include the proof here for completeness.
Lemma 4.1.
The graphs $S_{2}$ and $T_{1}$ do not have the *-C1P.
Proof.
Since the argument are the same for both graphs, we use the same notations for their vertices. Let $a^{\prime}$, $b^{\prime}$ and $c^{\prime}$ be the three leaves in $S_{2}$ (respectively $T_{1}$), $a$, $b$ and $c$ their respective neighbors and let $x$ the central vertex of $S_{2}$ (see Figure 5). Suppose that $S_{2}$ (respectively $T_{1}$) has the *-C1P. Then there exists an ordering $\sigma$ on the vertices such that for each vertex $u$, there exists $N_{u}\in\{N[u],N(u)\}$ (in particular, $N(u)\subseteq N_{u}\subseteq N[u]$), such that $N_{u}$ consists of consecutive vertices in the ordering $\sigma$. Observe that in both graphs, the sets $N_{a}$, $N_{b}$, and $N_{c}$ are pairwise intersecting. Since these sets correspond to consecutive sets in the ordering $\sigma$, one of those sets, w.l.o.g., $N_{a}$, is contained in the union of the others, here $N_{b}$ and $N_{c}$. However, $a^{\prime}$ belongs to $N(a)\subseteq N_{a}$ and does not belong to $N[b]\cup N[c]$, and, hence, does also not belong to $N_{b}\cup N_{c}$, a contradiction. ∎
Figure 5: Notation for the vertices of $S_{2}$ and $T_{1}$ in Lemma 4.1, with the open neighborhoods of $a$, $b$, and $c$.
We restate our main theorem of this section, confirming a conjecture from [BHR25]. See 1.6
Proof.
It follows directly from the definition of *-C1P that if $G$ is a graph that has the *-C1P and $H$ is an induced subgraph of $G$, then $H$ also has the *-C1P. Thus, by Lemma 4.1, graphs with *-C1P are $\{S_{2},T_{1}\}$-free. Therefore, they are also $\{S_{2},T_{2}\}$-free and, hence, have path eccentricity at most 1 by Theorem 1.4. ∎
5 Algorithmic Proof (Theorem 1.8)
In this section, we use the following notation regarding paths in graphs. Given a path $P$ of extremities $u,v$ and vertices $x,y$ adjacent respectively to $u,v$, we denote by $xPy$ the path obtained by adding the edges $xu$, $vy$ to $P$. Similarly, we denote by $xP$ and $Py$ the paths obtained by adding respectively only $xu$ or only $vy$ to $P$.
See 1.8
Proof.
Let $P$ be a longest path in $G$. Suppose that $P$ is not dominating. In such a case, we show how to modify $P$ to obtain a longest path $P^{\prime}$ that dominates more vertices than $P$. Let $u,v$ be the extremities of $P$.
Since $P$ does not dominate $G$ and $G$ is connected, there exists a vertex $x$ at distance $2$ from $P$. Let $y$ be a neighbor of $x$ adjacent to some vertex of $P$, and $w$ be a neighbor of $y$ in $P$ that is closest to $u$ along $P$. Observe that $y$ and $w$ have to exist, as $x$ is at distance $2$ from $P$.
Observe that neither $u$ nor $v$ has a neighbor outside of $P$, as $P$ is a longest path. Hence, $w$ is an internal vertex of $P$, implying in particular that $u$, $v$, and $w$ are three distinct vertices of $P$. Denote by $u^{\prime}$, $v^{\prime}$ the neighbor of $u$, resp. $v$ on $P$. Observe that $u^{\prime}\neq v$ and $v^{\prime}\neq u$. Let $P^{*}$ be the subgraph of $P$ obtained from deleting $u$ and $v$ from $P$. In case $y$ is adjacent to any of $u^{\prime},v^{\prime}$, we define $P^{\prime}$ as $xyP^{*}v$ or $uP^{*}yx$. In both cases, $P^{\prime}$ has a greater length than $P$, a contradiction. Therefore, $x$ and $y$ are non-adjacent to any of $u$, $v$, $u^{\prime}$, and $v^{\prime}$. Moreover, as $y$ is adjacent to $P$, the vertices $u^{\prime}$, $w$, and $v^{\prime}$ are pairwise distinct, implying that $|V(P)|\geqslant 5$.
Let $P^{1}$ and $P^{2}$ be the two paths resulting from deleting $u$, $v$, and $w$ from $P^{*}$, containing respectively $u^{\prime}$ and $v^{\prime}$. Suppose that $u^{\prime}v\in E(G)$, or $uv^{\prime}\in E(G)$, or $uv\in E(G)$. Then, the path $Q$ defined respectively by $xywP^{1}vP^{2}$, or $xywP^{1}uP^{2}$, or $xywP^{1}uvP^{2}$ is longer than $P$, a contradiction. Therefore, $\{u^{\prime}v,uv^{\prime},uv\}\cap E(G)=\emptyset$. Suppose that $u^{\prime}$ is non-adjacent to $v^{\prime}$. Then, the set $\{u,v,u^{\prime},v^{\prime}\}$ induces a $2P_{2}$ in $G$ and consequently $\{x,y,u,v,u^{\prime},v^{\prime}\}$ induces a $3P_{2}$, a contradiction. Therefore, $u^{\prime}v^{\prime}\in E(G)$. Let $P^{\prime}$ be the path $xywP^{1}P^{2}$. Note that the path $wP^{1}P^{2}$ has vertex set $V(P)\setminus\{u,v\}$, and, since neither $u$ nor $v$ has a neighbor outside of $P$, the path $wP^{1}P^{2}$ dominates as many vertices as $P$. Hence, $P^{\prime}$ has the same length as $P$, but it dominates more vertices.
The above argument shows that we can modify a given longest path $P$ at most $|V(G)|$ times until we obtain a longest path that is dominating.
It remains to discuss the running time in the case we start the process above with some arbitrary (not necessarily longest) path. Then, in each step of the argument above, we modify the current path $P$ to a path $P^{\prime}$ that is either longer than $P$ or has the same length but dominates more vertices. Therefore, we run the standard BFS algorithm checking whether path $P$ is already dominating. If not, it finds a vertex $x$ at distance $2$ from $P$ together with a connecting path back to $P$ and we proceed with the algorithm finding $P^{\prime}$ and repeat. BFS runs in time $\mathcal{O}(n+m)$. Our proof guarantees at most $n^{2}$ steps, as in each step we either increase the size of the dominated set by at least $1$, or we do not decrease it but we increase the length of path $P$. ∎
6 Concluding Remarks and Open Questions
Our main result, Theorem 1.4, shows that for every integer $k\geqslant 1$, each connected $\{S_{k},T_{k}\}$-free graph has path eccentricity less than $k$. This raises the question regarding the complexity of computing the path eccentricity of such a graph. While the problem is clearly polynomial for $k=1$, it turns out that the problem is NP-hard for all $k\geqslant 2$. This follows from the fact that the problem of determining whether a given graph has a Hamiltonian path is NP-complete in the class of split graphs [Mül96], which are known to be $2P_{2}$-free and, hence, $\{S_{2},T_{2}\}$-free. This, however, says nothing about the complexity of computing a path witnessing the upper bound on path eccentricity guaranteed by Theorem 1.4.
Question 6.1.
Is there an algorithmic version of Theorem 1.4? More precisely: for a fixed integer $k\geqslant 1$ and a connected $\{S_{k},T_{k}\}$-free graph $G$, can we find in polynomial time a path in $G$ with eccentricity less than $k$?
While 6.1 is known to have an affirmative answer for $k=1$ (see [She91, BDK00]), it is, to the best of our knowledge, open for all $k\geqslant 2$. In fact, even a weaker version of 6.1, finding a path of eccentricity less than $k$ in cases covered by Corollary 1.7, is open, except for the case of $3P_{2}$-free graphs, which is covered by Theorem 1.8.
Another result of this paper (Theorem 1.8) is that every connected $3P_{2}$-free graph $G$ has a longest path that is dominating. This result, along with the observation that there exist graphs that are not $3P_{2}$-free such that every connected induced subgraph has a longest path that is dominating (consider, for instance, paths and cycles), raise the question about a characterization of such graphs.
Question 6.2.
What are the graphs $G$ such that every connected induced subgraph of $G$ has a longest path that is dominating?
Camby and Schaudt showed in [CS16] that for every integer $k\geqslant 4$, every connected $P_{k}$-free graph $G$, and every minimum connected dominating set $S$ in $G$, the subgraph of $G$ induced by $S$ is either $P_{k-2}$-free or isomorphic to $P_{k-2}$. Can some similar properties regarding minimum connected dominating sets (at distance) be proved for $\{S_{k},T_{k}\}$-free graphs? In particular:
Question 6.3.
Is it true that for every integer $k\geqslant 2$, every connected $\{S_{k},T_{k}\}$-free graph has a minimum connected distance-$(k-1)$ dominating set inducing a $\{S_{1},T_{1}\}$-free graph?
Acknowledgements
The authors are grateful to Ulrich Pferschy for pointing out reference [ADG+24], which motivated the present research.